# Evaluating the World Model Implicit in a Generative Model

### v1

@@ 1 Introduction
We study whether a language model trained on sequences of directions recovers the true underlying map. This question relates to other state tracking problems studied in the language modeling literature. For example, Patel & Pavlick [24] show that larger LLMs ground spatial concepts like cardinal directions to locations in a grid world and generalize to various grid layouts. Additionally, Guan et al. [8] use LLMs to perform planning tasks from natural language descriptions. Our results suggest that LLMs can perform some of these tasks well (such as finding shortest paths between two points on a map) without having a coherent world model.
@@ 3.2 Evaluating world models
To implement the sequence compression metric, we randomly sample states (i.e., [intersection, destination] pairs) and two distinct traversals that arrive at each state. We then assess whether a model correctly admits the same suffixes for each prefix. We average over pairs of prefixes to report a score for each state and average over states to report a final score. To implement the sequence distinction metrics, we sample pairs of distinct states and a traversal that arrives at the state, comparing the model’s approximate Myhill-Nerode boundary to the true one. We average over pairs of prefixes to report a score for each pair of states, and average over 1000 randomly sampled state pairs to report a final scores. Both metrics depend on a threshold parameter $\epsilon$: a prefix is only sampled or accepted if the model’s assigned probability for each token is above $\epsilon$. Here, we consider $\epsilon=0.01$ for all models and metrics. We describe implementation details and provide parameter ablations in Appendix D.
@@ 3.3 Reconstructing implicit maps
Our evaluation metrics point to deficiencies in recovering world models. We now show that these metrics reveal underlying incoherence. In the maps setting, the state structure of the true world model is easy to interpret and visualize: it is defined by the map itself. We attempt to “reconstruct” the map implied by the sequences produced by each generative model.
@@
Figure 3 shows three reconstructed maps using sequences generated by the transformer trained on random walks. On the left is the reconstructed map on only sequences which are valid under the true world model. On the right is the reconstructed map using the transformer’s sequences. The transformer’s underlying world model is incoherent; it recovers streets whose orientations are physically impossible (e.g. labeled NW but facing east) and require flyovers above other streets.
@@ Acknowledgements
Keyon Vafa is supported by the Harvard Data Science Initiative. Justin Chen is supported by an NSF Graduate Research Fellowship under Grant No. 174530. Jon Kleinberg is supported in part by a Vannevar Bush Faculty Fellowship, a Simons Collaboration grant, and a grant from the MacArthur Foundation. We thank the Chicago Booth School of Business for generous support. We thank Foundry for providing the compute required to conduct this research. We also thank Juan Carlos Perdomo and Neekon Vafa for helpful comments and feedback.
@@ B.1 Algorithm
Our reconstruction procedure in Algorithm 1 assumes, when possible, that sequences output by the transformer which are consistent with the world model correspond to edges structure that the transformer has grasped. In a sense, we give the generative model the benefit of the doubt. However, sometimes the model errs, which means that in the transformer’s map, some edge must exist which does not exist in the world model. In this case, our reconstruction algorithm must choose a new edge to add. We add a new edge that maximizes the amount of steps we can take without again running into a model error. In other words, we add new edges only when we encounter a discrepancy between the transformer and the world model, and we choose the edge which greedily maximizes the number of succeeding steps which correspond to the world model.
@@ B.2 Maps
We include reconstructed maps built from $6400$ transformer-generated sequences for transformers trained on shortest paths, noisy shortest paths (modeling traffic), and random walks in Manhattan. We vary the constrained maximum degree between $4$ and $8$ (the true map has maximum degree $4$ and there are $8$ possible cardinal directions), and the maximum edge distance between $\nicefrac{{1}}{{2}}$ and $1$ mile.
@@
We note that the maps corresponding to the random walk model visually have more edges than, for instance, the maps built from shortest paths data. This is simply due to the fact that while the inputs to the reconstruction algorithm have the same number of sequences, the length of the sequences differ between the different generative models. The total number of directions contained in the shortest path sequences is $983182$, for noisy shortest paths is $1140487$, and for random walks is $1584549$.

### v2

@@ 1 Introduction
We study whether a language model trained on sequences of directions recovers the true underlying map. This question relates to other state tracking and navigation problems studied in the language modeling literature [27, 28]. For example, Patel & Pavlick [24] show that larger LLMs ground spatial concepts like cardinal directions to locations in a grid world and generalize to various grid layouts. Relatedly, Schumann & Riezler [26] demonstrate that transformer-based models can generate navigation instructions in language from underlying graphs. Additionally, Guan et al. [8] use LLMs to perform planning tasks from natural language descriptions. Our results suggest that LLMs can perform some of these tasks well (such as finding shortest paths between two points on a map) without having a coherent world model.
@@ 3.2 Evaluating world models
To implement the sequence compression metric, we randomly sample states (i.e., [intersection, destination] pairs) and two distinct traversals (i.e. prefixes) that arrive at each state. We then assess whether a model correctly admits the same suffixes for each prefix. We average over pairs of prefixes to report a score for each state and average over states to report a final score. To implement the sequence distinction metrics, we sample pairs of distinct states and traversals (i.e. prefixes) that arrive at each state, comparing the model’s approximate Myhill-Nerode boundary to the true one. We average over pairs of prefixes to report a score for each pair of states, and average over 1000 randomly sampled state pairs to report a final scores. Both metrics depend on a threshold parameter $\epsilon$: a prefix is only sampled or accepted if the model’s assigned probability for each token is above $\epsilon$. Here, we consider $\epsilon=0.01$ for all models and metrics. We describe implementation details and provide parameter ablations in Appendix D.
@@ 3.3 Reconstructing implicit maps
Our evaluation metrics point to deficiencies in recovering world models. We now show that these metrics reveal underlying incoherence. In the maps setting, the state structure of the true world model is easy to interpret and visualize: it is defined by the map itself. We attempt to “reconstruct” the map implied by sequences sampled from each generative model.
@@
Figure 3 shows three reconstructed maps using sequences generated by the transformer trained on random walks. The sequences underlying each map are generated by randomly sampling 6400 (origin, destination) pairs and then sampling the model’s traversal for each pair. On the left is the reconstructed map on only sequences which are valid under the true world model. On the right is the reconstructed map using the transformer’s sequences. The transformer’s underlying world model is incoherent; it recovers streets whose orientations are physically impossible (e.g. labeled NW but facing east) and require flyovers above other streets.
@@ Acknowledgements
Keyon Vafa is supported by the Harvard Data Science Initiative. Justin Chen is supported by an NSF Graduate Research Fellowship under Grant No. 174530. Jon Kleinberg is supported in part by a Vannevar Bush Faculty Fellowship, a Simons Collaboration grant, and a grant from the MacArthur Foundation. We thank the Chicago Booth School of Business for generous support. We thank Foundry [^4] https://www.mlfoundry.com/ for providing the compute required to conduct this research. We also thank Juan Carlos Perdomo and Neekon Vafa for helpful comments and feedback.
@@ B.1 Algorithm
Our reconstruction algorithm in Algorithm 1 takes a set of sequences from a generative model and attempts to reconstruct the underlying map implied by the sequences. The reconstructed map has the same set of vertices as the true world model (i.e. the intersections of Manhattan), and we visualize each intersection by placing them at their real-world latitude/longitude. The reconstruction algorithm thus attempts to recover the edges and edge directions implied by each set of sequences.
For each sequence that is valid under the true map, the algorithm adds the true edges and edge directions implied by the sequence. In a sense, this algorithm gives the generative model the benefit of the doubt. However, sometimes the model errs, i.e. it produces a sequence which is only valid by adding an additional edge to the true map. In this case, our reconstruction algorithm adds a new edge at the first invalid step of the traversal. There are usually multiple possible edges to add that would make the traversal valid; the algorithm adds the edge that maximizes the number of subsequent steps that would be valid. In other words, the algorithm adds new edges only when there is a discrepancy between the transformer and the world model, choosing the edge which greedily maximizes the number of subsequent steps that accord with the true world model. The direction label for each added edge is thus the token at the corresponding invalid step of the traversal.
@@ B.2 Maps
We include reconstructed maps built from $6400$ transformer-generated sequences for transformers trained on shortest paths, noisy shortest paths (modeling traffic), and random walks in Manhattan. Each sequence is generated by randomly sampling an (origin, destination) pair and then sampling the model’s traversal for each pair. Because the transformer never sees sequences of length more than 100 during training, we only sample pairs for which there exists a valid traversal in less than 100 moves. This distribution of (origin, destination) pairs still varies from the distributions used to train and evaluate each model, and the percent of traversals that are valid falls from >95% to 65-80%. We vary the constrained maximum degree between $4$ and $8$ (the true map has maximum degree $4$ and there are $8$ possible cardinal directions), and the maximum edge distance between $\nicefrac{{1}}{{2}}$ and $1$ mile.
Each map depicts the map of Manhattan implied by the graph reconstruction algorithm. Reconstructing maps involves adding edges between two intersections. We make sure the edges visually accord with the labels reconstructed by the algorithm. For example, if intersection $v_{1}$ is north of intersection $v_{2}$ in the true map but the reconstruction algorithm recovers an edge from $v_{2}$ to $v_{1}$ labeled “South”, we draw an edge that leaves $v_{2}$ facing south and loops back to $v_{1}$. In the zoomed-in images, edges belonging to the true map are in black and false edges added by the reconstruction algorithm are in red.
@@
We note that the maps corresponding to the random walk model visually have more edges than, for instance, the maps built from shortest paths data. This is due to the fact that while the inputs to the reconstruction algorithm have the same number of sequences, the length of the sequences differ between the different generative models. The total number of directions contained in the shortest path sequences is $983182$, for noisy shortest paths is $1140487$, and for random walks is $1584549$.